# An ordering for the strength of functional dependence

### v1

@@ 1.3 Main results
The following main result shows that the conditional convex order in Definition 1.2 satisfies all the proposed axioms for a global dependence ordering.
@@
For a closed set $A\subseteq[0,1]\,$, define $\mathcal{R}_{A}:=\{(Y,\mathbb{X})\colon\Omega\to\mathbb{R}^{1+p}\mid\overline{\mathsf{Ran}(F_{Y})}=A\,,\,p\in\mathbb{N}\}\,.$ Then, for $\mathcal{R}=\mathcal{R}_{A}$, the conditional convex order $\preccurlyeq_{ccx}$ in Definition 1.2 satisfies the Axioms 18.
@@
    The class $\mathcal{R}_{A}$ in Theorem 1.4 consists of all random vectors $(Y,\mathbb{X})$ that are at least two-dimensional and satisfy the marginal constraint with respect to the set $A$. For example, if $A=[0,1]$, then $\mathcal{R}_{A}$ includes all random vectors $(Y,\mathbb{X})$ such that $Y$ has a continuous distribution function; note that there is no assumption on the distribution or dimension of $\mathbb{X}$. The set $\mathcal{R}_{A}$ includes the extreme elements in conditional convex order, for example, the comonotonic random vector $(Y,Y)$ as maximal element and the independent random vector $(Y,c)$, $c\in\mathbb{R}$, as minimal element.
@@
    The conditional convex order is not symmetric, i.e., for bivariate random vectors $(Y,X)$ and $(Y^{\prime},X^{\prime})$, the relation $(Y,X)\preccurlyeq_{ccx}(Y^{\prime},X^{\prime})$ does not imply $(X,Y)\preccurlyeq_{ccx}(X^{\prime},Y^{\prime})$. This is consistent with the asymmetric notion of perfect dependence, where $Y$ being a function of $X$ does not imply that $X$ is a function of $Y$.
@@
    As a consequence of the information monotonicity and the characterization of conditional independence, the conditional convex order satisfies the data processing inequality $(Y,\mathbb{Z})\preccurlyeq_{ccx}(Y,\mathbb{X})$ whenever $Y$ and $\mathbb{Z}$ are conditionally independent given $\mathbb{X}\,.$ It follows that $(Y,\mathbb{h}(\mathbb{X}))\preccurlyeq_{ccx}(Y,\mathbb{X})$ for all measurable functions $\mathbb{h}$. Further, the conditional convex order is self-equitable in the sense that $(Y,\mathbb{h}(\mathbb{X}))=_{ccx}(Y,\mathbb{X})$ for all measurable functions $\mathbb{h}$ such that $Y$ and $\mathbb{X}$ are conditionally independent given $\mathbb{h}(\mathbb{X})\,.$
In the next theorem, we characterize the conditional convex order through a comparison of bivariate SI random vectors in concordance order $\leq_{c}$ (defined in (27)). To this end, consider for $(Y,\mathbb{X})$ the rearranged quantile transform $(0,1)^{2}\ni(t,u)\mapsto q_{Y|\mathbb{X}}^{\uparrow u}(t)\in\mathbb{R}$ defined in Equation (10) below. We will consider reduced random vectors of the form $(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)$, which are bivariate random vectors that have the same degree of dependence as $(Y,\mathbb{X})$ in conditional convex order. We refer to Proposition 2.7 for their basic properties. We denote by $U(0,1)$ the uniform distribution on $(0,1)$, assuming that the underlying probability space $(\Omega,\mathcal{A},P)$ is non-atomic and thus rich enough to ensure the existence of $U(0,1)$-distributed random variables.
@@
    $(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)\leq_{c}(F_{Y^{\prime}}\circ q_{Y^{\prime}|\mathbb{X}^{\prime}}^{\uparrow U}(V),U)\,.$
In particular, $(Y,\mathbb{X})=_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ is equivalent to $(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(F_{Y^{\prime}}\circ q_{Y^{\prime}|\mathbb{X}^{\prime}}^{\uparrow U}(V),U)\,.$
@@
    The concordance order requires equal marginal distributions; hence, we have
    $F_{Y}(q_{Y|\mathbb{X}}^{\uparrow U}(V))\stackrel{{\scriptstyle\mathrm{d}}}{{=}}F_{Y^{\prime}}(q_{Y^{\prime}|\mathbb{X}^{\prime}}^{\uparrow U}(V))$ in Theorem 1.6(ii). This condition is equivalent to the marginal constraint $\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}})}$ resulting from the definition of the conditional convex order; to see this, combine Proposition 2.7(i) and Lemma A.2.
@@
    If $F_{Y}$ is continuous, then $F_{Y}(q_{Y|\mathbb{X}}^{\uparrow U}(V))$ and $F_{Y^{\prime}}(q_{Y^{\prime}|\mathbb{X}^{\prime}}^{\uparrow U}(V))$ are uniform on $(0,1)$, and the distribution functions of the reduced random vectors in Theorem 1.6(ii) are SI copulas. Consequently, the conditional convex order can be reduced to a pointwise comparison of SI copulas. In particular, Theorem 1.6 strengthens Axiom 8 in the sense that the conditional convex order only depends on the bivariate SI copula obtained from the dimension reduction principle in Proposition 2.7.
The following theorem shows that the strength of functional dependence in the additive error model (1) decreases in $\sigma$ with respect to the conditional convex order. Hence, as a consequence of Theorem 4.1, Chatterjee’s rank correlation and related dependence measures are decreasing in the parameter $\sigma$ of the additive error model.
@@
    The proof of Theorem 1.8 is based on Theorem 2.2 which characterizes the conditional convex order by decreasing rearrangements of transformed conditional distribution functions. If $Y$ follows a convolution semigroup, the behavior of $\xi$ can be verified much easier using the data processing inequality for $\xi$ in [17]: In this case, $Y$ admits a representation $Y=X_{1}+\varepsilon_{1}+\varepsilon_{2}$ for independent random variables $X_{1},\varepsilon_{1},\varepsilon_{2}$; see e.g. [38, Theorem 14.50]. For $X_{2}:=X_{1}+\varepsilon_{1}$, it follows that $(Y\mid X_{2}=x)\sim\varepsilon_{2}+x$ and $(X_{1}\mid X_{2}=x)\sim x-\varepsilon_{1}$. Hence, $Y$ and $X_{1}$ are conditionally independent given $X_{2}$, which implies $\xi(Y,X_{1})\leq\xi(Y,X_{2})$. More generally, the data processing inequality for the conditional convex order yields $(Y,X_{1})\preccurlyeq_{ccx}(Y,X_{2})$; see Remark 1.5(d).
1.4 Related literature and our contribution
The growing need to evaluate the significance of features in large and complex datasets has, in recent years, motivated research on association measures satisfying various desirable axioms, as studied in [13, 37, 44, 49, 53]. Dependence measures that characterize independence and perfect functional dependence have been studied by Dette et al. [22] and Gamboa et al. [26], and recently in [5, 7, 8, 13, 15, 21, 29, 61, 62]. In their seminal work, Azadkia and Chatterjee [7] developed a nearest-neighbor-based estimator for the rank correlation $\xi$ in (2), and they established a powerful, model-free variable selection method. Their method relies on the fact that $\xi$ not only satisfies the zero-independence and max-functionality property but also characterizes conditional independence, fulfills information monotonicity, and satisfies several invariance properties.
While measures of association like the Pearson correlation, Spearman’s rho or Kendall’s tau have a clear interpretation, there is still lack in interpreting dependence measures like $\xi$. Closed-form expressions are only rarely available for some specific distributions and measures; see e.g. [2] for a closed-form expression of $\xi$ for the multivariate normal distribution. To better understand the properties and behaviors of such dependence measures, an ordering that underlies these measures is of fundamental importance. It is well known that measures of concordance, such as Kendall’s $\tau$, Spearman’s $\rho$, or Gini’s $\gamma$, are increasing in the concordance order [32, Section 2.12]. In contrast, for dependence measures like Chatterjee’s $\xi$ an open problem concerns the identification of a dependence ordering that underlies these measures. We resolve this problem and contribute to the literature on dependence measures as follows:
@@ 1.4 Related literature and our contribution
  • We propose Axioms 18 as natural and desirable axioms for a global dependence ordering that describes the strength of functional dependence of a random variable $Y$ on a random vector $\mathbb{X}$. As we show in Theorem 1.4, the conditional convex order introduced in Definition 1.2 satisfies all these axioms. To the best of our knowledge, there is no work that addresses all these axioms, and we are not aware of any dependence ordering that has all these properties. For an overview of dependence orderings, we refer to [32, 47, 57].
@@
  • In Section 2, we derive two characterizations of the conditional convex order. The first one relies on the Schur order used to compare the variability of conditional distribution functions; see Theorem 2.2. As an example, we verify $\preccurlyeq_{ccx}$ for bivariate Bernoulli distributions; see Proposition 2.5. For the second characterization, we introduce a dimension reduction principle to transform a random vector $(Y,\mathbb{X})$ to a bivariate SI random vector. Due to Proposition 2.7, this reduced random vector has the same strength of functional dependence in the sense of the conditional convex order. In our second main result, Theorem 1.6, we characterize $\preccurlyeq_{ccx}$ by comparing reduced random vectors in concordance order.
  • In Section 3, we provide sufficient conditions for verifying $\preccurlyeq_{ccx}$ in the additive error model (1) (Theorem 1.8) and for various copula-based models (Theorem 3.3). For the multivariate normal distribution, we obtain a notable characterization of the conditional convex order and show that multivariate normal random vectors are always comparable in $\preccurlyeq_{ccx}$ (Theorem 3.1).
  • As shown by Borgonovo et al. [13], convexity is essential for a dependence measure to satisfy information monotonicity and to characterize independence and perfect dependence. In Theorem 4.1 and Theorem 4.4, we use the conditional convex order to generate new dependence measures that are increasing in $\preccurlyeq_{ccx}$, possess the zero-independence and max-functionality property, characterize conditional independence, satisfy information monotonicity, and exhibit various useful invariance properties. We are not aware of related construction methods that yield dependence measures with all these properties.
@@
  • In Section 5, we compare the conditional convex order with related positive and global dependence orderings from the literature [22, 58, 59, 63, 68]. We show that none of the existing orderings satisfies Axioms 18, nor do they describe the behavior of Chatterjee’s rank correlation across various models.
@@
Throughout the paper, boldface symbols $\mathbb{x}$ and $\mathbb{X}$ are used for vectors and random vectors, respectively, and standard symbols $y$ and $Y$ for real numbers and real-valued random variables. All proofs are deferred to the Appendix B.
@@ 2 Characterizations of $\preccurlyeq_{ccx}$
In this section, we derive two characterizations of the conditional convex order. The first is based on the Schur order, which compares the variability of integrable functions via their decreasing rearrangements. The second characterization is formulated in terms of the concordance order applied to dimension-reduced bivariate SI random vectors.

### v2

@@ 1.3 Main results
The following main result shows that the conditional convex order in Definition 1.2 satisfies all the proposed axioms for a global dependence ordering. Therefore, we consider for a closed set $A\subseteq[0,1]$ the class
$\mathcal{R}_{A}:=\{(Y,\mathbb{X})\colon(\Omega,\mathcal{A},P)\to(\mathbb{R}^{1+p},\mathcal{B}(\mathbb{R}^{1+p}))\mid\overline{\mathsf{Ran}(F_{Y})}=A\,,\,p\in\mathbb{N}\}$ (6)
of random vectors.
@@
For $\mathcal{R}=\mathcal{R}_{A}$, the conditional convex order $\preccurlyeq_{ccx}$ in Definition 1.2 satisfies the Axioms 18.
@@
    The class $\mathcal{R}_{A}$ in Theorem 1.4 consists of all random vectors $(Y,\mathbb{X})$ that are at least two-dimensional and satisfy the marginal constraint with respect to the set $A$. For example, if $A=[0,1]$, then $\mathcal{R}_{A}$ includes all random vectors $(Y,\mathbb{X})$ such that $Y$ has a continuous distribution function; note that there is no assumption on the distribution or dimension of $\mathbb{X}$. The set $\mathcal{R}_{A}$ includes the extreme elements in $\preccurlyeq_{ccx}$, for example, the comonotone random vector $(Y,Y)$ as maximal element and the independent random vector $(Y,c)$, $c\in\mathbb{R}$, as minimal element.
@@
    The conditional convex order is not symmetric, i.e., for bivariate random vectors $(Y,X)$ and $(Y^{\prime},X^{\prime})$, the relation $(Y,X)\preccurlyeq_{ccx}(Y^{\prime},X^{\prime})$ does not imply $(X,Y)\preccurlyeq_{ccx}(X^{\prime},Y^{\prime})$. This is in line with the asymmetric notion of perfect dependence, where $Y$ being a function of $X$ does not imply that $X$ is a function of $Y$.
@@
    As a consequence of information monotonicity and characterization of conditional independence, the conditional convex order satisfies the data processing inequality $(Y,\mathbb{Z})\preccurlyeq_{ccx}(Y,\mathbb{X})$ whenever $Y$ and $\mathbb{Z}$ are conditionally independent given $\mathbb{X}\,.$ It follows that $(Y,\mathbb{h}(\mathbb{X}))\preccurlyeq_{ccx}(Y,\mathbb{X})$ for all measurable functions $\mathbb{h}$. Further, the conditional convex order is self-equitable in the sense that $(Y,\mathbb{h}(\mathbb{X}))=_{ccx}(Y,\mathbb{X})$ for all measurable functions $\mathbb{h}$ such that $Y$ and $\mathbb{X}$ are conditionally independent given $\mathbb{h}(\mathbb{X})\,.$
In the next theorem, our second main result, we give two characterizations of the conditional convex order—both in terms of bivariate SI random vectors. These random vector are compared with respect to the concordance order $\leq_{c}$, the positive dependence ordering mentioned above; see also Section 5.1. The first characterization relies on conditional comonotonicity—a dependence concept that underlies Wasserstein correlations. The second characterization is based on a dimension reduction principle that we study in detail in Section 2.2. Regarding the notation, we consider for $(Y,\mathbb{X})$ the rearranged quantile transform $(0,1)^{2}\ni(t,u)\mapsto q_{Y;\mathbb{X}}^{\uparrow u}(t)$ defined in Equation (13) below. Then, for independent $U,V\sim U(0,1)$, the bivariate random vector $(F_{Y}\circ q_{Y;\mathbb{X}}^{\uparrow U}(V),U)$ is equivalent to $(Y,\mathbb{X})$ in conditional convex order, that is,
$(F_{Y}\circ q_{Y;\mathbb{X}}^{\uparrow U}(V),U)=_{ccx}(Y,\mathbb{X});$ (7)
see Theorem 2.7. We refer to such a bivariate random vector as reduced random vector. As we study in Section 4, each of the equivalent versions in the following theorem generates a class of dependence measures that are $\preccurlyeq_{ccx}$-increasing and inherit all their fundamental properties from the conditional convex order.
@@
    $(F_{Y}\circ F_{Y|\mathbb{X}}^{-1}(V),V)\geq_{c}(F_{Y^{\prime}}\circ F_{Y^{\prime}|\mathbb{X}^{\prime}}^{-1}(V),V)$ with $V$ independent of $\mathbb{X}$ and $\mathbb{X}^{\prime}$,
  3. (iii)
    $(F_{Y}\circ q_{Y;\mathbb{X}}^{\uparrow U}(V),U)\leq_{c}(F_{Y^{\prime}}\circ q_{Y^{\prime};\mathbb{X}^{\prime}}^{\uparrow U}(V),U)\,.$
In particular, we have $(Y,\mathbb{X})=_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ if and only if $(F_{Y}\circ F_{Y|\mathbb{X}}^{-1}(V),V)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(F_{Y^{\prime}}\circ F_{Y^{\prime}|\mathbb{X}^{\prime}}^{-1}(V),V)$ if and only if $(F_{Y}\circ q_{Y;\mathbb{X}}^{\uparrow U}(V),U)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(F_{Y^{\prime}}\circ q_{Y^{\prime};\mathbb{X}^{\prime}}^{\uparrow U}(V),U)$.
@@
    Recall from (4) that the conditional convex order underlies Chatterjee’s rank correlation. In Theorem 4.1, we study classes of related dependence measures that are $\preccurlyeq_{ccx}$-increasing and constructed via conditional distribution functions.
@@
    By Theorem 1.6ii, $\preccurlyeq_{ccx}$ is characterized via a comparison of bivariate random vectors in concordance order. These random vectors are conditionally comonotone and SI. Conditional comonotonicity follows from the representation $P(F_{Y|\mathbb{X}}^{-1}(V)\leq y,V\leq y^{\prime})=\int_{\mathbb{R}^{p}}\min\{F_{Y|\mathbb{X}=\mathbb{x}}(y),y^{\prime}\}\mathsf{\,d}P^{\mathbb{X}}(\mathbb{x})$. Note that concavity of $u\mapsto\min\{u,y\}$ explains the reversed inequality sign in ii. Wasserstein correlations [68] are based on conditional comonotonicity and, as we show in Theorem 4.4, they are $\preccurlyeq_{ccx}$-increasing. While $\preccurlyeq_{ccx}$ is defined via distribution/survival functions, the characterization in ii is a quantile-based version. Independence of $(F_{Y|\mathbb{X}}^{-1}(V),V)$ characterizes perfect dependence of $Y$ on $\mathbb{X}$. Conversely, comonotonicity of $(F_{Y|\mathbb{X}}^{-1}(V),V)$ corresponds to independence of $\mathbb{X}$ and $Y$.
  3. (iii)
    Rearranged dependence measures [62] are constructed via measures of concordance and reduced random vectors; see (34). As a consequence of Theorem 1.6iii, such measures are $\preccurlyeq_{ccx}$-increasing; see Theorem 4.6. The reduced random vectors $(F_{Y}\circ q_{Y;\mathbb{X}}^{\uparrow U}(V),U)$ are SI. They are constructed via the dimension-reduction principle in Section 2.2, and, notably, they are $\preccurlyeq_{ccx}$-equivalent to $(Y,\mathbb{X})$; see (7). Independence/Comonotonicity of $(q_{Y;\mathbb{X}}^{\uparrow U}(V),U)$ characterizes independence/perfect dependence of $Y$ and/on $\mathbb{X}$.
  4. (iv)
    The concordance order requires equal marginal distributions. Consequently, we have $F_{Y}(F_{Y|\mathbb{X}}^{-1}(V))\stackrel{{\scriptstyle\mathrm{d}}}{{=}}F_{Y^{\prime}}(F_{Y^{\prime}|\mathbb{X}^{\prime}}^{-1}(V))$ and $F_{Y}(q_{Y;\mathbb{X}}^{\uparrow U}(V))\stackrel{{\scriptstyle\mathrm{d}}}{{=}}F_{Y^{\prime}}(q_{Y^{\prime};\mathbb{X}^{\prime}}^{\uparrow U}(V))$ in Theorem 1.6ii and iii. Either condition is equivalent to the marginal constraint $\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}})}$ resulting from the definition of the conditional convex order.
The next theorem, which constitutes our third main result, provides new insights into the behavior of dependence measures like $\xi$ beyond the boundary values 0 and $1$. While their construction is based on the zero-independence and max-functionality, these measures typically satisfy further useful properties such as information monotonicity; see [12]. However, as we discuss in Remark 1.9 below, information monotonicity alone does not explain monotonicity of dependence measures, for instance, in the noise level of additive error models. In Theorem 1.8, we show that the additive error model (1) is decreasing in the parameter $\sigma$ with respect to the conditional convex order, and hence so are all $\preccurlyeq_{ccx}$-increasing dependence measures.
@@
    Consider $Y_{\sigma}:=X+\sigma\varepsilon$ where $X$ and $\varepsilon$ are standard normal and independent. Then, for $0<\sigma^{\prime}<\sigma^{\prime\prime}$, we have $(Y_{\sigma^{\prime}},X)\succcurlyeq_{ccx}(Y_{\sigma^{\prime\prime}},X)$ by Theorem 1.8 and, obviously, $(Y_{\sigma^{\prime}},X)\not=_{ccx}(Y_{\sigma^{\prime\prime}},X)$. Since the $\sigma$-algebras generated by $(Y_{\sigma^{\prime}},X)$ and $(Y_{\sigma^{\prime\prime}},X)$ coincide, there is no information gain for $Y_{\sigma}$ in the sense of Axiom 5 when decreasing the parameter $\sigma$. Thus, information monotonicity is not sufficient for a comparison in $\preccurlyeq_{ccx}$.
1.4 Our contribution to the literature and structure of the paper
@@
We propose Axioms 18 as natural and desirable axioms for a global dependence ordering that describes the strength of functional dependence of a random variable $Y$ on a random vector $\mathbb{X}$. As shown in Theorem 1.4, our first main result, the conditional convex order introduced in Definition 1.2 satisfies all these axioms. To the best of our knowledge, there is no work that addresses all these axioms, and we are not aware of any dependence ordering that satisfies all these fundamental properties. We refer to [12, 37, 43, 50, 55] for various axioms at the level of measures of association, and to [32, 45, 58] for an overview of dependence orderings.
@@
In Theorem 1.6, our second main result, we present two characterizations of $\preccurlyeq_{ccx}$ based on the bivariate concordance order. The first characterization in Theorem 1.6ii relies on conditionally comonotone random vectors. The second characterization in Theorem 1.6iii is based on a dimension-reduction principle using the concept of rearrangements, which we study in detail in Section 2. More precisely, we use the Hardy-Littlewood-Polya theorem to rearrange conditional distribution functions and to characterize $\preccurlyeq_{ccx}$ in terms of the Schur order; see Theorem 2.2. This characterization is particularly useful for verifying $\preccurlyeq_{ccx}$ in several models. As a first example, we verify $\preccurlyeq_{ccx}$ for bivariate Bernoulli distributions in Proposition 2.5. The dimension reduction principle is formalized in Theorem 2.7 where we construct, for a given random vector $(Y,\mathbb{X})$, a $\preccurlyeq_{ccx}$-equivalent bivariate SI random vector. In Example 2.9, we illustrate the dimension-reduction procedure for the multivariate normal distribution.
Section 3 builds on the dimension reduction procedure and provides sufficient conditions for verifying $\preccurlyeq_{ccx}$ under stochastic monotonicity assumptions. In particular, we verify $\preccurlyeq_{ccx}$ for copula-based models in Theorem 3.3. This theorem is also used to prove Theorem 1.8 on additive error models. Somewhat surprisingly, Theorem 3.1 shows that multivariate normal distributions are always comparable in conditional convex order, and the strength of functional dependence is solely encoded in the Schur complements of the covariance matrices.
In Section 4, we show that the conditional convex order generates several classes of dependence measures recently studied in the statistics literature. Due to Theorem 4.1, $\preccurlyeq_{ccx}$ underlies Chatterjee’s rank correlation and related dependence measures constructed via conditional distribution functions. As we show in Theorem 4.4, $\preccurlyeq_{ccx}$ also underlies Wasserstein correlations [68]. In Theorem 4.6, we prove that rearranged dependence measures [62] are $\preccurlyeq_{ccx}$-increasing. Similar results hold true for the integrated $R^{2}$ [8] (see Remark 4.3b) and the kernel partial correlation [29] (see Remark 3.2b).
@@
In Section 5, we compare the conditional convex order with related positive and global dependence orderings from the literature [22, 59, 60]. In particular, we show that none of the existing orderings satisfies Axioms 18, nor do they describe the behavior of Chatterjee’s rank correlation and related dependence measure across various models.
@@
Some useful properties of generalized inverses, convex order, and transformations of random vectors are given in Appendix A. All proofs are deferred to Appendix B. Throughout the paper, boldface symbols $\mathbb{x}$ and $\mathbb{X}$ are used for vectors and random vectors, respectively, and standard symbols $y$ and $Y$ for real numbers and real-valued random variables.
@@ 2 Characterizations of $\preccurlyeq_{ccx}$
Recall Theorem 1.6 where two characterizations of the conditional convex order are given. In Theorem 2.2 below, we establish a third characterization of $\preccurlyeq_{ccx}$ in terms of the Schur order, which is based on rearrangements and can be verified in several settings. The link to the $\preccurlyeq_{ccx}$-order is provided by the Hardy–Littlewood–Pólya theorem (Lemma 2.1), which characterizes the (standard) convex order in terms of the Schur order. The dimension-reduction procedure in Section 2.2 relies on rearrangements and is of central importance for our paper.